Three kids in a parking lot launch a rocket that rises into air along a 380-m long arc in 40s. Determine its average speed.


1
Expert's answer
2021-03-08T08:30:36-0500
"v=\dfrac{d}{t}=\dfrac{380\ m}{40\ s}=9.5\ \dfrac{m}{s}."